A satellite is in circular orbit at an altitude of 1500 km above the surface of a nonrotating planet with an orbital speed of 3.9 km/s. The minimum speed needed to escape from the surface of the planet is 9.6 km/s, and G = 6.67 × 10-11 N · m2/kg2. The orbital period of the satellite is closest to

54 min.37 min.49 min.43 min.60 min.
Answer #1

60 min is the correct answer.

given

h = 1500 km = 1.5*10^6 m

vo = 3.9 km/s = 3900 m/s

ve = 9.6 km/s = 9600 m/s


let M is the mass of planet and R is the radius of planet


we know, orbital speed, vo = sqrt(G*M/(R+h))

escape speed, ve = sqrt(2*G*M/R)


ve/vo = sqrt(2*(R+h)/R)

(ve/vo)^2 = 2*(R + h)/R

(9600/3900)^2 = 2*(R + 1.5*10^6)/R

on solving the above equation we get

R = 7.39*10^5 m


radius of orbit, r = R + h

= 7.39*10^5 + 1.5*10^6

= 2.239*10^6 m

orbital periode, T = 2*pi*r/vo

= 2*pi*2.239*10^6/3900

= 3607 s

= 3607/(60)

= 60.1 min <<<<<--------------ANswer